OpenStudy (anonymous):

Use the Law of Sines to find the indicated angle θ. Assume C = 63°, b = 56.3, c = 80.1. (Round your answer to one decimal place.)

OpenStudy (dumbcow):

im assuming theta is angle B ?

OpenStudy (dumbcow):

sin B/b = sinC/c

OpenStudy (anonymous):

=> sinB = b sin C/c = 56.3 * sin 63/ 80.1 =

OpenStudy (anonymous):

56.3 * sin 63 degrees/ 80.1 = 0.626263013 ??

OpenStudy (anonymous):

B = arcsin ( .63) =38 degree

OpenStudy (callisto):

like dumbcow said (sin B)/b = (sinC)/c sinB = b (sinC ) /c = 56.3 x (sin 63) /80.1 = 0.626263012 that's the value of sin B only, B = sin^-1 (0.626263012) = 38.77 degrees